a.) Calculate the molarity of 0.150 mol of Na2S in 1.10 L of solution?

b.) calculate the molarity of 32.9g of MgS in 763 mL of solution?

c.) If 4.82 g of CuNO3 is dissolved in water to make a 0.830 M solution, what is the volume of the solution in milliliters?

d.) A 5.17 g sample of a laboratory solution contains 1.97 g of acid. What is the concentration of the solution as a mass percentage?


1
Expert's answer
2021-02-22T06:10:04-0500

a) molarity = moles / volume

Molarity= 0.150/1.10 = 0.136M

b) moles of MgS = 32.9/56.38 = 0.583 moles

Volume = 763ml = 0.763L

Molarity = moles / volume

Molarity = 0.583/0.763 = 0.764M

c) moles of CuNO3 = 4.82/125.55 = 0.033mole

Molarity = moles/volume

0.830 = 0.033/volume

Volume = 0.033/0.830 = 0.039L = 39 ml

d) %Mass = (1.97/5.17)×100 = 38.15%

%Mass = "\frac{5.17-1.97}{5.17}" ×100 = 61.85%